what is the acid- test ratio? Give at least 3 examples
Answer #1

The acid-test ratio uses a firm's balance sheet data as an indicator of whether it has sufficient short-term assets to cover its short-term liabilities. This metric is more useful in certain situations than the current ratio, also known as the working capital ratio, since it ignores assets such as inventory, which may be difficult to quickly liquidate.

The acid-test ratio is also commonly known as the quick ratio.

Key points -

1. The acid-test, or quick ratio, compares a company's most short-term assets to its most short-term liabilities to see if a company has enough cash to pay its immediate liabilities, such as short-term debt.

2. The acid test ratio disregards current assets that are difficult to liquidate quickly such as inventory.

3. The acid-test ratio may not give a reliable picture of a firm's financial condition if the company has accounts receivable that take longer than usual to collect or current liabilities that are due but have no immediate payment needed.

The Formula for the Acid-Test Ratio

Acid Test=

Cash + Marketable Securities+​ AR / Current liabilities

Where AR = accounts receivable


OR

QR = current assets - Inventory - prepaid / current liabilities

Example 1 -

Cash: $10,000
Accounts Receivable: $5,000
Inventory: $5,000
Stock Investments: $1,000
Prepaid taxes: $500
Current Liabilities: $15,000

QR = 10000 + 5000 + 1000 / 15000

= 1.067

Example 2

Inventory: $5,000
Prepaid taxes: $500
Total Current Assets: $21,500
Current Liabilities: $15,000

QR = CURRENT ASSET - inventory - prepaid tax / current liabilities

= 21500 - 5000 - 500 / 15000

= 1.067

Example 3

Cash and cash equivalents 21,120

Short-term investments 20,481

Receivables 16,849

Inventories 2,349

Deferred income taxes 5,546

Other current assets 23,033

Total current assets 89,378

Total current liabilities 80,610

Qr = 21120 + 20481 + 16849 / 80610

= 0.75